# Factorization in Finitely-Presented Monoids

### v1

@@ Abstract
We study arithmetic properties of factorizations of elements into products of generators, in monoids given with explicit presentations. After relating and comparing this perspective to the more usual approach of factoring into products of atoms, we explore how the relations in the presentation of a monoid affect factorization. In the process, we construct a large class of non-commutative fully elastic monoids. We also show that any finitely-presented cancellative normalizing monoid satisfies the Structure Theorem for Unions. Examples are constructed to demonstrate the sharpness of our results, and exhibit unusual factorization behavior.
@@ 1 Introduction
For a long time Factorization Theory was mainly focused on commutative integral domains and commutative cancellative monoids, and the relevant building blocks were atoms (i.e., non-units that cannot be expressed as products of non-units). In the last decade the scope has spread rapidly to new algebraic fields, and now includes commutative rings with zero-divisors, as well as non-commutative rings and monoids (with or without idempotents). This necessitated new building blocks, beyond atoms (see [41, 29, 16, 13, 26] for recent surveys).
@@
Transfer homomorphisms are a key tool in Factorization Theory. They preserve systems of factorization length sets, and they can provide a bridge between objects having quite different algebraic properties. Deep work has revealed that there are transfer homomorphisms from various classes of non-commutative prime rings to commutative monoids [42, 4, 40], from non-integrally closed domains to integrally closed domains [37], and more [5]. Nevertheless, such statements are far from being true in general. There are no length-preserving homomorphisms from Weyl algebras to commutative monoids [6, Lemma 8.1], or from general power monoids to cancellative monoids [19, Proposition 4.12]. There are finitely-generated right-cancellative monoids, presented by just one relation, whose systems of length sets are distinct from the system of length sets of any commutative finitely-generated cancellative monoid [28, Corollary 4.4]. Every finitely-generated cancellative monoid, that is commutative, has bounded factorization (i.e., all length sets are finite) and satisfies the Structure Theorem for Length Sets. Non-commutative finitely-presented cancellative monoids need not even have bounded factorization [6, Section 7]. It is well-known and easy to see that commutative noetherian domains have bounded factorization, but it is an open problem whether this remains true for non-commutative noetherian prime rings. These results demonstrate, not surprisingly, that arithmetic investigations of non-commutative monoids and rings need new approaches and techniques, beyond what was used in the commutative setting. Furthermore, the loss of commutativity, in classes such as finitely-generated cancellative monoids or noetherian rings, may change arithmetic properties substantially.
@@
In this note we study factorizations in non-commutative monoids presented by generators and relations. This line of inquiry started in [28], with atoms as building blocks. Indeed, in all the prior literature on factorizations there is only a small number of papers where factorization invariants are not defined via atoms (e.g., [30, 10]), though more recently, adaptations of factorization building blocks to other settings include radical ideals, valuation elements, and quarks (e.g., [9, 21, 13]). In this paper, instead of considering factorizations of elements into products of atoms, we use generators, since they provide a natural factorization basis in the presence of an explicit presentation. We reformulate the usual arithmetic concepts of Factorization Theory, such as length and elasticity, in terms of generators (Section 2), translate various standard results about elasticity and the Structure Theorem for Unions to that setting (using abstract results of Tringali [44] about subadditive collections of natural numbers), and then conduct a systematic study.
@@
In Section 3 we connect factorization in terms of generators to factorization in terms of atoms, and discuss the consequences of our framework. In particular, generators and atoms are very closely related, and for a reduced monoid (i.e., one having no non-identity units) with an irredundant generating set, being atomic is equivalent to the generators and atoms coinciding (Proposition 3). Then, in Section 4, we explore the manner in which the relations in the presentation of a monoid affect factorization. This aspect was largely neglected in the past, even in the setting of commutative and cancellative monoids (though some steps in that direction have been taken by Chapman, García-Sánchez, Philipp, et al., see [11, 35, 36]). Among other observations, we show that in a monoid presented by one relation, the factorization length sets form arithmetic progressions (Proposition 5). Moreover, any monoid with a free generator and accepted elasticity is fully elastic (Theorem 9).
@@
In Section 6 we give explicit constructions of finitely-presented cancellative monoids that have various pathological factorization behaviors, to demonstrate the sharpness of our results from Section 5. More specifically, in Proposition 18 we exhibit a one-relation cancellative bounded factorization monoid that does not have accepted elasticity, but satisfies the Structure Theorem for Unions. Then, in Proposition 21, we construct a two-relation cancellative bounded factorization monoid that does not satisfy the Structure Theorem for Unions. There are few monoids in the literature, of any flavor, that fail to satisfy that theorem, with the first appearing in [18].
@@ 2.1 General Terminology
Clearly, every cancellative or acyclic monoid is unit-cancellative, but not conversely. A cancellative non-commutative monoid need not be acyclic.
@@ 2.3 Factorizations
We say that $M$ is a bounded factorization, or BF, monoid, respectively half-factorial, if $\mathsf{L}_{M}(a)$ is finite, respectively $\mathsf{L}_{M}(a)=\{|a|\}$, for all $a\in\langle X\rangle$.
@@ 2.4 Structure Theorem for Unions
The three statements above follow from [44, Theorem 2.20], [44, Corollary 2.24], and [44, Theorem 2.27], respectively. ∎
@@
Since $X\neq\emptyset$, necessarily $\rho(M)\neq 0$, and so $\mathcal{L}(M)$ satisfies the hypotheses of [44, Proposition 2.12]. The equivalence of the above statements now follows from that result. ∎
3 Atoms vs. Generators
@@
Considering factorizations into generators, rather than atoms, has the advantage of being applicable to arbitrary monoids, even if they have no atoms at all (e.g., antimatter domains [3, 7]). Of course, as with atoms, the resulting arithmetic indicators are not always informative. For example, suppose that $M:=\langle X\mid R\rangle$ is such that $ab=_{M}1$, for some $a,b\in\langle X\rangle\setminus\{1\}$. Then for all $c\in\langle X\rangle$ and $n\in\mathbb{N}$, we have
@@
and so $\mathsf{L}_{M}(c)$ contains an infinite arithmetic progression with difference $|ab|$. In this situation the system of lengths of $M$ sheds relatively little light on the factorization properties of $M$.
One possible remedy to the specific problem above would be to consider only elements of $M$ that can be factored into products of generators $x\in X$ with the property that $axb\neq_{M}1$, for all $a,b\in\langle X\rangle$. This would be analogous to the standard approach, for commutative non-atomic monoids, of considering factorizations only for elements that can be written as products of atoms (since no product of atoms in a commutative monoid can be a unit). However, doing so would not rule out other similar problems that arise, for example, in semigroups with nontrivial idempotents.
Another option, which we deploy in Section 5, would be to foreclose all issues of this sort by assuming that the monoid is BF. The next lemma explores the consequences of that assumption. In particular, it translates to our setting the fact that, for atomic monoids, being BF implies being unit-cancellative (see [6, Lemma 2.2 and Theorem 2.3]).
@@
As noted in Section 2.3, the various factorization-related notions we have defined depend on the specific presentation of a given monoid. To take an extreme example, the trivial monoid $\{1\}$ is BF with respect to the standard presentation $\langle\emptyset\rangle$, but not with respect to the presentation $\langle x\mid(1,x)\rangle$. This is yet another consequence of the more general framework, and we tame it in many of our results by choosing generating sets with desirable features, such as being irredundant. In particular, for any monoid, the traditional atom-based approach can be reconstructed within our framework, by focusing on factorizations in the submonoid generated by the atoms.
@@ 4 Basic Observations
Since $a=_{M}b$ and $a\neq b$, there are $m\in\mathbb{N}^{+}$ and $c_{0},\dots,c_{m}\in\langle X\rangle$ such that $c_{0}=b$, $c_{m}=a$, and each $c_{i}$ is connected to $c_{i+1}$ via an elementary $R$-transition. That is, for each $i<m$, there exist $d_{i},g_{i},g_{i}^{\prime},h_{i}\in\langle X\rangle$ satisfying $(g_{i},g_{i}^{\prime})\in\{(e,f),(f,e)\}$, $c_{i}=d_{i}g_{i}h_{i}$, and $c_{i+1}=d_{i}g_{i}^{\prime}h_{i}$. (See [31, Proposition 1.5.9] for more details.) Thus $||c_{i+1}|-|c_{i}||=d$ for each $i$, and $\Delta(\mathsf{L}_{M}(a))=\{d\}$. Moreover, given that $|a|=\min(\mathsf{L}_{M}(a))$, and hence $|a|\leq|b|$, we have
@@
While one-relation monoids are well-behaved, in the above sense, they can exhibit some interesting factorization behaviors none-the-less. Specifically, we show in Proposition 18 that such a monoid need not have accepted elasticity, even when the elasticity is finite. Additionally, in Lemma 19 we construct a family of one-relation monoids $M$ having the unusual property that the differences $\rho_{k}(M)-\rho_{k-1}(M)$ are finite but unbounded. We also note that a (commutative) one-relation monoid need not be fully elastic (cf. Theorem 9 below). Specifically, if $n_{1},n_{2}\in\mathbb{N}$ satisfy $1<n_{1}<n_{2}$ and $\gcd(n_{1},n_{2})=1$, then the submonoid $M$ of $(\mathbb{N},+)$ generated by $\{n_{1},n_{2}\}=\mathcal{A}(M)$ is not fully elastic (see [27, Theorem 5.5] for more details). More generally speaking, one-relation monoids have historically served as a rich source for interesting questions–see, e.g., [33].
@@
(1) Let $a,b\in\langle X\rangle$ be such that $a=_{M}b$. If $a=b$, then $|a|-|b|=0\in N$, and so we may assume that $a\neq b$. Then there must exist $m\in\mathbb{N}^{+}$ and $c_{0},\dots,c_{m}\in\langle X\rangle$ such that $c_{0}=b$, $c_{m}=a$, and each $c_{i}$ is connected to $c_{i+1}$ via an elementary $R$-transition [31, Proposition 1.5.9]. Given that $R$ is symmetric, this means that for each $i<m$ there exist $d_{i},e_{i},e_{i}^{\prime},f_{i}\in\langle X\rangle$, such that $(e_{i},e_{i}^{\prime})\in R$, $c_{i}=d_{i}e_{i}f_{i}$, and $c_{i+1}=d_{i}e_{i}^{\prime}f_{i}$. Then
@@
According to [18, Proposition 2.9], $\min(\Delta(\mathcal{L}))=\gcd(\Delta(\mathcal{L}))$, for any directed collection $\mathcal{L}$ of subsets of $\mathbb{N}$. Since $N\neq\{0\}$ implies that $X\neq\emptyset$, the system of lengths $\mathcal{L}(M)$ is indeed directed (see Section 2.3), and so we conclude that $d=\gcd(\Delta(M))$. ∎
The following is a generalization of [23, Proposition 6].
@@
$(3)\Rightarrow(4)$ This follows from the fact that if $k\in\mathbb{N}$ and $a\in\langle X\rangle$ is such that $|a|=k$, then
@@
Next, given that $M$ has accepted elasticity, for all $a\in\langle X\rangle\setminus\{1\}$ and $k\in\mathbb{N}^{+}$, we have

### v2

@@ Abstract
We study arithmetic properties of factorizations of elements into products of generators, in monoids given with explicit presentations. After relating and comparing this perspective to the more usual approach of factoring into products of atoms, as well as other more recent alternatives, we explore how the relations in the presentation of a monoid affect factorization. In the process, we construct a large class of non-commutative fully elastic monoids. We also show that any finitely-presented cancellative normalizing monoid satisfies the Structure Theorem for Unions. Examples are constructed to demonstrate the sharpness of our results, and exhibit unusual factorization behavior.
@@ 1 Introduction
For a long time Factorization Theory was mainly focused on commutative integral domains and commutative cancellative monoids, and the relevant building blocks were atoms (i.e., non-units that cannot be expressed as products of non-units). In the last decade the scope has spread rapidly to new algebraic fields, and now includes commutative rings with zero-divisors, as well as non-commutative rings and monoids (with or without idempotents). This necessitated new building blocks, beyond atoms (see [52, 37, 22, 17, 34] for recent surveys; further specific references are given below).
@@
Transfer homomorphisms are a key tool in Factorization Theory. They preserve systems of factorization length sets, and they can provide a bridge between objects having quite different algebraic properties. Deep work has revealed that there are transfer homomorphisms from various classes of non-commutative prime rings to commutative monoids [51, 4, 53], from non-integrally closed domains to integrally closed domains [47], and more [5]. Nevertheless, such statements are far from being true in general. There are no length-preserving homomorphisms from Weyl algebras to commutative monoids [6, Lemma 8.1], or from general power monoids to cancellative monoids [25, Proposition 4.12]. There are finitely-generated right-cancellative monoids, presented by just one relation, whose systems of length sets are distinct from the system of length sets of any commutative finitely-generated cancellative monoid [36, Corollary 4.4]. Every finitely-generated cancellative monoid, that is commutative, has bounded factorization (i.e., all length sets are finite) and satisfies the Structure Theorem for Length Sets. Non-commutative finitely-presented cancellative monoids need not even have bounded factorization [6, Section 7]. It is well-known and easy to see that commutative noetherian domains have bounded factorization, but it is an open problem whether this remains true for non-commutative noetherian prime rings. These results demonstrate, not surprisingly, that arithmetic investigations of non-commutative monoids and rings need new approaches and techniques, beyond what was used in the commutative setting. Furthermore, the loss of commutativity, in classes such as finitely-generated cancellative monoids or noetherian rings, may change arithmetic properties substantially.
Until quite recently, in the literature on factorizations contained only a small number of papers where factorization invariants are not defined via atoms (e.g., [38, 13]). However, in an effort to generalize the theory and overcome some of the aforementioned hurdles, factorization building blocks have been subsequently adopted to other settings, such as radical ideals, valuation elements, and others (e.g., [11, 27, 42, 12, 40]). Moreover, a general preorder-based theory of factorization has been rapidly developing, where a wide array of elements can be used as building blocks (see [55, 18, 19, 20, 21]).
@@
This note is devoted to factorizations in not-necessarily-commutative monoids presented by generators and relations–a line of inquiry started in [36], with atoms as building blocks. (For earlier contributions in the commutative cancellative setting see [14, 28]). Here, however, instead of considering factorizations of elements into products of atoms (or other aforementioned alternatives), we use generators, since they provide a natural factorization basis in the presence of an explicit presentation. We reformulate the usual arithmetic concepts of Factorization Theory, such as length and elasticity, in terms of generators (Section 2), translate various standard results about elasticity and the Structure Theorem for Unions to that setting (using abstract results of Tringali [54] about subadditive collections of natural numbers), and then conduct a systematic study.
@@
In Section 3 we connect factorization in terms of generators to factorization in terms of atoms, as well as other types of building blocks appearing in recent literature, and discuss the consequences of our framework. In particular, generators and atoms are very closely related, and for a reduced monoid (i.e., one having no non-identity units) with an irredundant generating set, being atomic is equivalent to the generators and atoms coinciding (Proposition 3). Then, in Section 4, we explore the manner in which the relations in the presentation of a monoid affect factorization. This aspect was largely neglected in the past, even in the setting of commutative and cancellative monoids (though some steps in that direction have been taken by Chapman, García-Sánchez, Philipp, et al.–see [15, 45, 46]). Among other observations, we show that in a monoid presented by one relation, the factorization length sets form arithmetic progressions (Proposition 6). Moreover, any monoid with a free generator and accepted elasticity is fully elastic (Theorem 10).
@@
In Section 6 we give explicit constructions of finitely-presented cancellative monoids that have various pathological factorization behaviors, to demonstrate the sharpness of our results from Section 5. More specifically, in Proposition 19 we exhibit a one-relation cancellative bounded-factorization monoid that does not have accepted elasticity, but satisfies the Structure Theorem for Unions. Then, in Proposition 22, we construct a two-relation cancellative bounded-factorization monoid that does not satisfy the Structure Theorem for Unions. There are few monoids in the literature, of any flavor, that fail to satisfy that theorem, with the first appearing in [24].
@@ 2.1 General Terminology
Clearly, every cancellative or acyclic monoid is unit-cancellative, but not conversely. A cancellative non-commutative monoid need not be acyclic. The concept of unit-cancellativity was first introduced in the commutative setting in [50], although under a different name, and independently in [24]. We would be remiss not to mention that the monoid-theoretic notion of unit-cancellativity is reminiscent of Bouvier’s notion of présimplifiable for commutative rings (see [9, 10]). The concept of acyclicity was introduced in [55].
@@ 2.3 Factorizations
We say that $M$ is a bounded-factorization, or BF, monoid, respectively is half-factorial, if $\mathsf{L}_{M}(a)$ is finite, respectively $\mathsf{L}_{M}(a)=\{|a|\}$, for all $a\in\langle X\rangle$.
@@ 2.4 Structure Theorem for Unions
The three statements above follow from [54, Theorem 2.20], [54, Corollary 2.24], and [54, Theorem 2.27], respectively. ∎
@@
Since $X\neq\emptyset$, necessarily $\rho(M)\neq 0$, and so $\mathcal{L}(M)$ satisfies the hypotheses of [54, Proposition 2.12]. The equivalence of the above statements now follows from that result. ∎
3 Generators vs. Atoms and Quarks
@@
Considering factorizations into generators, rather than atoms, has the advantage of being applicable to arbitrary monoids, even if they have no atoms at all (e.g., antimatter domains [3, 7]). Moreover, generator-based factorization can be viewed as a general platform for exploring factorization into building blocks of various sorts, as we describe more explicitly below.
Remark 4.
Let $M$ be a monoid, let $X\subseteq M\setminus\{1\}$, and let $N$ be the submonoid of $M$ generated by $X$. Then $N\cong\langle X\mid\ker(\phi)\rangle$, where $\phi:\langle X\rangle\to N$ is the homomorphism induced by sending each element of $X$ to the corresponding copy in $N$, and $\phi(1)=1$. Hence taking $X$ to be the set of building blocks for a theory of factorization in $M$ (e.g., primes, atoms, irreducible elements), one can reformulate that theory in terms of generators, upon restricting to the submonoid $N$ consisting of the elements of $M$ that can be factored into products of the relevant building blocks.
The essence of the idea of restricting to the submonoid generated by factorization building blocks is commonly used in, for example, the Factorization Theory of commutative non-atomic monoids, where factorizations are considered only for elements that can be written as products of atoms. Likewise, using presentations consisting of factorization building blocks is not new. Specifically, the factorization monoid, which is the free commutative monoid generated by the atoms of a given commutative monoid, is a commonly-used tool in the factorization literature (see, e.g., [45, 46, 4, 24, 37, 33]). Rather, our framework is a natural development of the earlier approaches, that is particularly well-suited to working in the setting where one typically constructs examples, namely monoids with explicit presentations.
A drawback of factorization into generators, as noted in Section 2.3, is that the various factorization-related notions we have defined depend on the specific presentation of a given monoid. To take an extreme example, the trivial monoid $\{1\}$ is BF with respect to the standard presentation $\langle\emptyset\rangle$, but not with respect to the presentation $\langle x\mid(1,x)\rangle$. Moreover, as with atoms, and other types of factorization building blocks, the generator-based arithmetic indicators are not always informative. For example, suppose that $M:=\langle X\mid R\rangle$ is such that $ab=_{M}1$, for some $a,b\in\langle X\rangle\setminus\{1\}$. Then for all $c\in\langle X\rangle$ and $n\in\mathbb{N}$, we have
@@
and so $\mathsf{L}_{M}(c)$ contains an infinite arithmetic progression with difference $|ab|$. In this situation the system of lengths of $M$ sheds relatively little light on the factorization properties of $M$. Similar problems arise, for example, in semigroups with nontrivial idempotents.
We tame problematic behavior of this sort in many of our results by choosing generating sets with desirable features, such as being irredundant. A more conventional approach is to impose some sort of regularity condition on the monoid, rather than the generators. We do precisely that in Section 5, where we assume, in most of the results, that our monoids are BF. The next lemma, versions of which can be found in various other settings (e.g., [25, Corollary 2.29], [18, Example 5.4], [6, Lemma 2.2 and Theorem 2.3]), explains some of the consequences of that assumption.
@@
We conclude this section by relating our setting to the preorder-based theory of factorization developed recently by Cossu and Tringali in [55, 18, 19, 20, 21]. Letting $M$ be a monoid equipped with a preorder $\preceq$ (i.e., reflexive transitive relation), one can define factorization building blocks with respect to that preorder, such as $\preceq$-irreducible elements, $\preceq$-atoms, $\preceq$-quarks, and $\preceq$-prime elements (see [55, Definition 3.6]). For example, $a\in M$ is a $\preceq$-quark if $b\prec a$ implies that $b\preceq 1\preceq b$, for all $b\in M$, and it is not the case that $a\preceq 1\preceq a$. Various results are then proved about factoring elements into products of the above types of building blocks, depending on properties of the preorder.
By Remark 4, this theory can be situated within our framework (though that may not necessarily be enlightening), by taking the $\preceq$-irreducible elements, or other types of elements mentioned above, as the generators in the presentation of an appropriate submonoid. On the other hand, our approach can be rephrased, in many cases, using the Cossu/Tringali terminology. More specifically, let $M:=\langle X\mid R\rangle$, and for all $a,b\in M$ write $a\preceq b$ if and only if $\min(\mathsf{L}(a))\leq\min(\mathsf{L}(b))$. Then $\preceq$ is a preorder on $M$ (which is a pullback, in the sense of [55, Example 3.3(2)], of the usual order on $\mathbb{N}$), and the $\preceq$-quarks are precisely the $a\in M$ such that $a=_{M}x$, for some $x\in X$. Thus, given a reasonably well-behaved generating set $X$, the generators of $M$ can be identified with its $\preceq$-quarks.
@@ 4 Basic Observations
Since $a=_{M}b$ and $a\neq b$, there are $m\in\mathbb{N}^{+}$ and $c_{0},\dots,c_{m}\in\langle X\rangle$ such that $c_{0}=b$, $c_{m}=a$, and each $c_{i}$ is connected to $c_{i+1}$ via an elementary $R$-transition. That is, for each $i<m$, there exist $d_{i},g_{i},g_{i}^{\prime},h_{i}\in\langle X\rangle$ satisfying $(g_{i},g_{i}^{\prime})\in\{(e,f),(f,e)\}$, $c_{i}=d_{i}g_{i}h_{i}$, and $c_{i+1}=d_{i}g_{i}^{\prime}h_{i}$. (See [39, Proposition 1.5.9] for more details.) Thus $||c_{i+1}|-|c_{i}||=d$ for each $i$, and $\Delta(\mathsf{L}_{M}(a))=\{d\}$. Moreover, given that $|a|=\min(\mathsf{L}_{M}(a))$, and hence $|a|\leq|b|$, we have
@@
While one-relation monoids are well-behaved, in the above sense, they can exhibit some interesting factorization behaviors none-the-less. Specifically, we show in Proposition 19 that such a monoid need not have accepted elasticity, even when the elasticity is finite. Additionally, in Lemma 20 we construct a family of one-relation monoids $M$ having the unusual property that the differences $\rho_{k}(M)-\rho_{k-1}(M)$ are finite but unbounded. We also note that a (commutative) one-relation monoid need not be fully elastic (cf. Theorem 10 below). Specifically, if $n_{1},n_{2}\in\mathbb{N}$ satisfy $1<n_{1}<n_{2}$ and $\gcd(n_{1},n_{2})=1$, then the submonoid $M$ of $(\mathbb{N},+)$ generated by $\{n_{1},n_{2}\}=\mathcal{A}(M)$ is not fully elastic (see [35, Theorem 5.5] for more details). More generally speaking, one-relation monoids have historically served as a rich source of interesting questions–see, e.g., [43].
@@
(1) Let $a,b\in\langle X\rangle$ be such that $a=_{M}b$. If $a=b$, then $|a|-|b|=0\in N$, and so we may assume that $a\neq b$. Then there must exist $m\in\mathbb{N}^{+}$ and $c_{0},\dots,c_{m}\in\langle X\rangle$ such that $c_{0}=b$, $c_{m}=a$, and each $c_{i}$ is connected to $c_{i+1}$ via an elementary $R$-transition [39, Proposition 1.5.9]. Given that $R$ is symmetric, this means that for each $i<m$ there exist $d_{i},e_{i},e_{i}^{\prime},f_{i}\in\langle X\rangle$, such that $(e_{i},e_{i}^{\prime})\in R$, $c_{i}=d_{i}e_{i}f_{i}$, and $c_{i+1}=d_{i}e_{i}^{\prime}f_{i}$. Then
@@
According to [24, Proposition 2.9], $\min(\Delta(\mathcal{L}))=\gcd(\Delta(\mathcal{L}))$, for any directed collection $\mathcal{L}$ of subsets of $\mathbb{N}$. Since $N\neq\{0\}$ implies that $X\neq\emptyset$, the system of lengths $\mathcal{L}(M)$ is indeed directed (see Section 2.3), and so we conclude that $d=\gcd(\Delta(M))$. ∎
The following is a generalization of [31, Proposition 6].
@@
$(3)\Rightarrow(4)$ This follows from the fact that if $k\in\mathbb{N}$ and $a\in\langle X\rangle$ are such that $|a|=k$, then
@@
Next, for all $a\in\langle X\rangle\setminus\{1\}$ and $k\in\mathbb{N}^{+}$, we have